# thanks: Corresponding author

### v1

@@ F.2 ISCO ($\epsilon=1$)
we note that, if the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ is momentarily ignored, the zeros of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ coincide with those of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{0}$, since both are controlled entirely by $\mathrm{d}M/\mathrm{d}r_{h}$. Because $\partial_{M}\Phi_{1}<0$ and $\mathrm{d}M/\mathrm{d}r_{h}$ increases monotonically, ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ is strictly decreasing.
Restoring the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}<0$ can only shift the zero of $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{1}$further to the left. Since the zero of $\mathrm{d}M/\mathrm{d}r_{h}$ already lies to the left of $r_{h,s}$ (as shown earlier), its additional leftward shift necessarily keeps it on the small-black-hole side.
Therefore, under the isothermal constraint, although $r_{\mathrm{c}}(r_{h})$ of the four-dimensional RN-AdS black hole is globally non-monotonic, it remains locally monotonic within the phase-transition region, thereby preserving the discontinuous jump across the transition.
Figure 9: Plots of $p(r_{h0})$ (red) and $p(r_{h,\max})$ (blue) as functions of $\tilde{T}_{0}$. At $\tilde{T}_{0}=\sqrt{2}/2$, one has $p(r_{h})>0$ for all $r_{h}$.
Figure 10: Behaviors of $P(r_{h})$ and $\mathrm{d}M/\mathrm{d}r_{h}$ for three representative values of $\tilde{T}_{0}$.

### v2

@@ Appendix G Monotonicity of $r_{\rm c}(r_{h})$ in the Phase Transition Region
we note that, if the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ is temporarily neglected, the zeros of $(\mathrm{d}M/\mathrm{d}r_{h})\,\partial_{M}\Phi_{1}$ coincide with those of $(\mathrm{d}M/\mathrm{d}r_{h})\,\partial_{M}\Phi_{0}$, since both are determined solely by $\mathrm{d}M/\mathrm{d}r_{h}$. Moreover, $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ is an increasing function of $r_{h}$ in a neighborhood of the zero. Restoring $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ then adds a positive contribution $(1/\Gamma_{1})\,\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}>0$, which lifts the curve and shifts the zero of $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ to smaller $r_{h}$. Consequently, the extremum of $r_{\rm isco}(r_{h})$ lies to the left of that of $r_{\rm ps}(r_{h})$, and hence also to the left of $r_{h,s}$, so the phase transition again occurs on the increasing branch of $r_{\rm isco}(r_{h})$.
Therefore, under the isothermal constraint, although $r_{\mathrm{c}}(r_{h})$ of the four-dimensional RN–AdS black hole is globally non-monotonic, it remains locally monotonic within the phase transition region, thereby preserving the discontinuous jump across the phase transition.
Figure 9: Plots of $\tilde{r}_{h,e}$, $\tilde{r}_{h,s}$, and $\tilde{r}_{h,l}$ as functions of $\tilde{T}_{0}$. The curves $\tilde{r}_{h,s}$ and $\tilde{r}_{h,l}$ intersect at point A, where $r_{h,s}=r_{h,l}=r_{h,c}$.